# From Source to Target: Leveraging Transfer Learning for Predictive Process Monitoring in Organizations

### v1

@@ 3 Practical relevance of transfer learning for predictive process monitoring in organizations
Transferring and utilizing knowledge from a source context S to a target context T is particularly interesting in connection with business processes. In organizations, IT systems are in place to support the execution of business processes to generate value. The widespread use of ERP systems offering similar modules for purchasing, sales and distribution, production, warehouse and inventory management, human resource management, or financial and cost accounting makes business processes comparable across organizations. Alongside ERP systems, the use of function-specific software, such as Salesforce for customer relationship management or ServiceNow for IT service management, leads to a standardization of processes. In addition to IT systems that offer a common ground for the execution of business processes, organizations can follow industry best practices and standards, for example, by the International Organization for Standardization (ISO)  [^1] https://www.iso.org (Accessed 23 July 2025) or the International Electrotechnical Commission (IEC)  [^2] https://www.iec.ch/homepage (Accessed 23 July 2025). In IT, standards include ISO/IEC 27001 for information security management systems or the Information Technology Infrastructure Library (ITIL)  [^3] https://www.axelos.com/certifications/itil-service-management/ (Accessed 23 July 2025) for IT service management. In manufacturing, ISO 9001 for quality management or Six Sigma for process improvement are common. Besides the voluntary adherence to industry standards, the activities and processes of an organization are also governed by rules and regulations. Prime examples of such regulations and directives are General Data Protection Regulation (GDPR)  [^4] https://gdpr.eu (Accessed 23 July 2025) and the recently introduced Corporate Sustainability Reporting Directive (CSRD)  [^5] https://finance.ec.europa.eu (Accessed 23 July 2025). Similarly, organizations must follow certain accounting standards based on their location, such as Generally Accepted Accounting Principles (GAAP)  [^6] https://fasab.gov/accounting-standards/ (Accessed 23 July 2025) or International Financial Reporting Standards (IFRS)  [^7] https://www.ifrs.org (Accessed 23 July 2025). Lastly, organizations have similar objectives based on which their processes are designed. For example, an organization can strive for cost or time reduction, improved efficiency, increased product or service quality, or customer satisfaction in their process executions. The use of comparable software tools, adherence to regulations, and (industry) standards, as well as common objectives, make different business processes within and across organizations increasingly similar to each other.
@@
This offers opportunities to transfer process knowledge within one organization (see Figure 9). Intra-organizational TL for PPM allows an organization to transfer knowledge across its processes. This might be especially useful when considering an organization operating with multiple subsidiaries in various locations. Such conglomerates often maintain multiple ERP systems for their subsidiaries based on industry or location. For example, an organization C has PPM with event data from its ERP system already in use in a source subsidiary $C_{S}$. To harness existing process knowledge in subsidiary $C_{S}$, a target subsidiary $C_{T}$ can utilize subsidiary $C_{S}$’s prediction models for its process data. This becomes inherently convenient when laws and regulations inhibit the collection of sufficient historical process data for subsidiary $C_{T}$ to train its own prediction models. In addition, intra-organizational TL for PPM offers the possibility for an organization to go beyond PPM pilot projects and provide process owners with the tools to implement PPM by themselves. Pre-trained prediction models for particular processes can be accessed by process owners to apply PPM without the need for costly resources and extensive model training, almost like an internal model-as-a-service offering.
@@
While transferring process knowledge through TL within one organization mainly benefits the same organization, TL for PPM can also be applied across organizational boundaries (see Figure 9). We coin this inter-organizational TL for PPM. For example, a target organization T is interested in the application of PPM and is planning to launch a pilot project to investigate its use for one of its support processes, incident management. Organization T is logging all information related to solving incidents in its service desk using ServiceNow, following ITIL while adhering to ISO/IEC 27001. With PPM, organization T aims to reduce the resolution time. To keep resources and costs low and to be able to launch a proof-of-concept relatively fast, organization T can use a transferred prediction model, trained with event data of a similar incident management process of source organization S instead of training a prediction model from scratch. Another use case for inter-organizational TL for PPM represents the acquisition of organization T by organization S. Some processes of organization S are already supported by PPM. As part of the merger, organization S wants to integrate organization T and its ERP system in its PPM efforts. To enable smooth integration, the prediction models used for processes in organization S can be transferred to similar processes in organization T.
@@
The increasing similarities of business processes create an application area for intra- and inter-organizational TL in PPM. As inter-organizational TL in PPM aims to transfer process knowledge across organizations, the domains (i.e., the feature space and marginal probability distribution of event data) of the source and target contexts tend to differ more than in intra-organizational TL in PPM, in which knowledge is transferred between business processes within the same organization. This greater difference in the source and the target domains in both application cases can be attributed to, for example, organization-specific standards and regulations for performing the same type of business process.
@@ 4 Related work
Related TL-based PPM techniques can generally be categorized into two groups: i) PPM techniques that transfer resources within one business process in an organization (i.e., $\pazocal{D}_{S}$ and $\pazocal{D}_{T}$ represent event data on the same business process) and ii) PPM techniques that transfer resources between similar business processes within one organization or across different organizations (i.e., $\pazocal{D}_{S}$ and $\pazocal{D}_{T}$ represent event data on similar business processes within the same or different organizations).
@@
Other PPM techniques in the first group aim to transfer parameters between DL models to improve prediction performance. Mehdiyev et al. (2020) propose a technique for the next activity prediction, in which two autoencoder models are trained in a semi-supervised manner on available event data ($\pazocal{D}_{S}$). The parameters from the hidden layers of these models are then transferred to a prediction model for the next activity prediction, which is subsequently fine-tuned using the same event data ($\pazocal{D}_{T}$). The work of Folino et al. (2019) suggests a technique for outcome prediction with limited labeled event data. An LSTM-encoder model is trained on event data ($\pazocal{D}_{S}$) for next activity prediction, as no extra labels are required for this target. The last hidden state of the encoder model’s final LSTM layer is then transferred to a second model for outcome prediction. This model is then fine-tuned using the same event data ($\pazocal{D}_{T}$) for which a process-related outcome label is available.
@@
Besides parameter-based TL, Cabrera et al. (2022) propose a PPM technique that leverages feature-based TL. This is achieved by using various BERT models – pre-trained on a huge corpus of text data ($\pazocal{D}_{S}$) – to encode the text information available in event data ($\pazocal{D}_{T}$) and adding the encoded information to the input of a multi-tasking LSTM model. This LSTM model addresses the next activity prediction and the next timestamp prediction. In addition, the work suggests that fine-tuning the pre-trained BERT models on available event data using the next activity prediction task enhances the LSTM model’s prediction performance for next activities and timestamps as it adapts the representation of the BERT models to the process control-flow.
@@
Some PPM techniques in the first group also perform both parameter- and feature-based TL. Pfeiffer et al. (2021) propose a PPM technique combining Gramian angular fields and representation learning. The Gramian angular fields are applied to encode the event data as 2D images, while representation learning is performed to obtain a generic process representation from the encoded event data ($\pazocal{D}_{S}$). This representation is obtained using a two-step approach. First, a CNN model is trained to map the encoded input onto trace variants. Second, the pre-trained CNN model is extended with a fully-connected layer and heads for various prediction tasks, such as the next activity and next timestamp prediction, and then fine-tuned on the encoded event data. Subsequently, all heads except the desired one are removed, and the model is fine-tuned again on the encoded input ($\pazocal{D}_{T}$). By learning a model capturing a generic process representation and aligning it to various prediction tasks, feature- and parameter-based TL is performed within one model. In a similar approach, Chen et al. (2022) propose to learn a generic representation of traces by pre-training a BERT on a set of unlabeled trace data ($\pazocal{D}_{S}$) using a custom masked activity modeling task. The pre-trained BERT model is then fine-tuned on another set of labeled trace data ($\pazocal{D}_{S}$) for next activity prediction or outcome prediction. As in the approach of Pfeiffer et al. (2021), one model capturing a generic process representation is learned and aligned to prediction tasks. Therefore, this approach also performs feature- and parameter-based TL. Pasquadibisceglie et al. (2023) suggest using TL for handling concept drift in PPM. Specifically, once a drift is detected within event data ($\pazocal{D}_{S}$), the parameters of both an LSTM model for the next activity prediction and a word2vec model for activity encoding are updated with current event data ($\pazocal{D}_{T}$). This improves the prediction performance of the LSTM model and enables the word2vec to deal with newly occurring activities. However, even though these PPM techniques incorporate parameter- and feature-based TL, they are not designed to work with different business processes, reflected by general differences in event attribute values and distributions in the source and the target domains, which is necessary for intra- or inter-organizational TL settings.
@@
Pfeiffer et al. (2021) Inductive, homogeneous Parameter-based, feature-based CNN model with task-specific layer ACT, TS, RES, DA Transfer process knowledge from a model capturing a generic business process representation for various single prediction tasks to prediction models for single tasks to address them more effectively Integer encoding, min-max scaling, Gramian angular field transformation Extraction of one time-related feature (duration since start), min-max normalization, Gramian angular field transformation RTP, NAP, OP Helpdesk, BPIC12, BPIC13, BPIC17, BPIC20, MobIS
Chen et al. (2022) Inductive, homogeneous Parameter-based, feature-based Multi-tasking, BERT-based transformer model ACT Transfer process knowledge from a generic business process representation for multiple tasks to prediction models for single tasks to address them more effectively Embedding layer, positional encoding - NAP, OP Helpdesk, BPIC12, Sepsis
@@
Cao et al. (2022) Transductive, homogeneous Parameter-based MLP model ACT, TS Transfer process knowledge from a prediction model pre-trained on event data (previous transition division) to another prediction model fine-tuned on event data (current transition division), to address a prediction task more effectively One-hot encoding, autoencoder Extraction of eight time-related features (e.g., month, execution time, elapsed time), min-max normalization, autoencoder RTP Four synthetic event logs, BPIC12, BPIC17,
@@
Cabrera et al. (2022) Inductive, heterogeneous Feature-based Multi-tasking LSTM model ACT, TS, text Transfer process knowledge from a generic embedding model pre-trained on a large corpus of text data to an embedding model fine-tuned on real-life event data, to address prediction tasks with a successive prediction model more effectively Embedding model Extraction of three time-related features (time passed since previous event, time within the day, time within the week), min-max normalization NAP, NTP BPIC16
@@
Nur et al. (2024) Transductive, heterogeneous Parameter-based, feature-based Transformer model with time-related feature block ACT, TS Transfer process knowledge from a prediction model pre-trained on real-life event data to another prediction model fine-tuned on real-life event data, to address a prediction task more effectively Embedding layer, positional encoding Extraction of two time-related features (elapsed time, lagged time) RTP Helpdesk, BPIC20, BPIC12, BPIC11
@@
Furthermore, two PPM techniques combine feature-based and parameter-based TL. Rizk et al. (2023) pre-train an LSTM model on synthetic event data ($\pazocal{D}_{S}$), transfer parameters of this model to a new one, and fine-tune the new model using other synthetic event data ($\pazocal{D}_{T}$) to solve the same prediction task (i.e., next activity prediction). The first layer of the LSTM model is an embedding layer, which enables matching the feature space of $\pazocal{D}_{S}$ to that of $\pazocal{D}_{T}$. The same experiment is performed for the event data of two different real-life event logs. Nur et al. (2024) suggest a technique with a transformer-based DL model for remaining time prediction. They emphasize the use of TL to address the challenge of limited event data. Their approach involved training a base model on a large amount of event data ($\pazocal{D}_{S}$) and fine-tuning it on a smaller amount of event data ($\pazocal{D}_{T}$) to enhance prediction performance for remaining process times. The framework integrates trace encoding, a transferable component block leveraging multi-head self-attention to capture semantic and temporal relationships, and time-related features. Although both techniques incorporate the ideas of feature-based and parameter-based TL in their models, they require a fine-tuning step on event data in the target context. However, this is difficult in intra- or inter-organizational TL settings, where event data are not or very limitedly available in the target context.
Table 1 summarizes the works identified, which use TL for PPM, and shows how they differ from our proposed technique in terms of dimensions, such as transfer goal, different process domains, and fine-tuning in target not necessary.
@@ 5 Transfer learning-based technique for predictive process monitoring
This section describes the design of our proposed technique for TL in PPM. The technique is structured into three phases, as shown in Figure 11. In the first phase, initial model building on source, an event log from the source context is loaded, preprocessed, and a DNN-based base model is built and evaluated on preprocessed event data. In the second phase, transfer resources from source to target, relevant resources from the source context (e.g., pre-trained models) are transferred to the target context. If an event log from the target context is available, the base model from the source context can be applied and evaluated in an intermediate step before being used for (ongoing) process instances in the target context. In the third phase, online application of model to target, an ongoing process instance of the target context is received. This instance is then pre-processed using the transferred resources, and the transferred base model is applied to the pre-processed data. Finally, operational support is provided based on the base model’s predictions for process improvement. In the following sections, the three phases of our proposed TL-based technique for PPM are described.

### v2

@@ 3 Practical relevance of transfer learning for predictive process monitoring in organizations
Transferring and utilizing knowledge from a source context S to a target context T is particularly interesting in connection with business processes. In organizations, IT systems are in place to support the execution of business processes to create value. The widespread use of ERP systems offering similar modules for purchasing, sales and distribution, production, warehouse and inventory management, human resource management, or financial and cost accounting makes business processes comparable across organizations. Alongside ERP systems, the use of function-specific software, such as Salesforce for customer relationship management or ServiceNow for IT service management, leads to a standardization of processes. In addition to IT systems that offer a common ground for executing business processes, organizations can follow industry best practices and standards. For example, they can follow those of the International Organization for Standardization (ISO)  [^1] https://www.iso.org (Accessed 23 July 2025) or the International Electrotechnical Commission (IEC)  [^2] https://www.iec.ch/homepage (Accessed 23 July 2025). In IT, standards include ISO/IEC 27001 for information security management systems or the Information Technology Infrastructure Library (ITIL)  [^3] https://www.axelos.com/certifications/itil-service-management/ (Accessed 23 July 2025) for IT service management. In manufacturing, ISO 9001 for quality management or Six Sigma for process improvement are common. Besides the voluntary adherence to industry standards, the activities and processes of an organization are also governed by rules and regulations. Prime examples of such regulations and directives are the General Data Protection Regulation (GDPR)  [^4] https://gdpr.eu (Accessed 23 July 2025) and the recently introduced Corporate Sustainability Reporting Directive (CSRD)  [^5] https://finance.ec.europa.eu (Accessed 23 July 2025). Similarly, organizations must follow certain accounting standards according to their location, such as the Generally Accepted Accounting Principles (GAAP)  [^6] https://fasab.gov/accounting-standards/ (Accessed 23 July 2025) or the International Financial Reporting Standards (IFRS)  [^7] https://www.ifrs.org (Accessed 23 July 2025). Lastly, organizations have similar objectives based on which their processes are designed. For example, an organization can strive for cost or time reduction, improved efficiency, increased product or service quality, or customer satisfaction in their process executions. The use of comparable software tools, adherence to regulations and (industry) standards, and common objectives make different business processes within and across organizations increasingly similar to each other.
@@
This offers opportunities to transfer process knowledge within one organization (see Figure 9). Intra-organizational TL for PPM allows an organization to transfer knowledge across its processes. This might be especially useful when considering an organization operating with multiple subsidiaries in various locations. Such conglomerates often maintain multiple ERP systems for their subsidiaries based on industry or location. For example, an organization C has PPM with event data from its ERP system already in use in a source subsidiary $C_{S}$. To harness existing process knowledge in subsidiary $C_{S}$, a target subsidiary $C_{T}$ can utilize subsidiary $C_{S}$’s prediction models for its process data. This becomes inherently convenient when laws and regulations inhibit the collection of sufficient historical process data for subsidiary $C_{T}$ to train its own prediction models. In addition, intra-organizational TL for PPM offers the possibility for an organization to go beyond PPM pilot projects and provide process owners with the tools to implement PPM by themselves. Pre-trained prediction models for certain process types can be accessed by process owners to apply PPM without the need for costly resources and extensive model training, almost like an internal model-as-a-service offering.
@@
While transferring process knowledge through TL within one organization mainly benefits the same organization, TL for PPM can also be applied across organizational boundaries (see Figure 9). We coin this inter-organizational TL for PPM. For example, a target organization T is interested in the application of PPM and plans to launch a pilot project to investigate its use for one of its support processes, incident management. Organization T logs all information related to solving incidents in its service desk using ServiceNow, following ITIL while adhering to ISO/IEC 27001. With PPM, organization T aims to reduce the resolution time. To keep resources and costs low and to be able to launch a proof-of-concept relatively fast, organization T can use a transferred prediction model, trained with event data of a similar incident management process of source organization S instead of training a prediction model from scratch. The acquisition of organization T by organization S represents another use case for inter-organizational TL for PPM. Some processes of organization S are already supported by PPM. As part of the merger, organization S wants to integrate organization T and its ERP system into its PPM efforts. To enable smooth integration, the prediction models used for processes in organization S can be transferred to similar processes in organization T.
@@
The increasing similarities of business processes create an application area for intra- and inter-organizational TL in PPM. As inter-organizational TL for PPM aims to transfer process knowledge across organizations, the domains (i.e., the feature space and marginal probability distribution of event data) of the source and target contexts tend to differ more than in intra-organizational TL for PPM, in which knowledge is transferred between business processes within the same organization. This greater difference between the source and target domains in an inter-organizational setting can be attributed to, for example, organization-specific standards and regulations for performing the same type of business process.
@@ 4 Related work
Related TL-based PPM techniques can generally be categorized into two groups: i) PPM techniques that transfer resources within one business process in an organization (i.e., $\pazocal{D}_{S}$ and $\pazocal{D}_{T}$ represent event data on the same business process), and ii) PPM techniques that transfer resources between similar business processes within one organization or across different organizations (i.e., $\pazocal{D}_{S}$ and $\pazocal{D}_{T}$ represent event data on similar business processes within the same or different organizations).
@@
Other PPM techniques in the first group aim to transfer parameters between DL models to improve prediction performance. Mehdiyev et al. (2020) propose a technique for the next activity prediction, in which two autoencoder models are trained in a semi-supervised manner on available event data ($\pazocal{D}_{S}$). The parameters from the hidden layers of these models are then transferred to a prediction model for the next activity prediction, which is subsequently fine-tuned using the same event data ($\pazocal{D}_{T}$). The work of Folino et al. (2019) suggests a technique for outcome prediction with limited labeled event data. An LSTM-encoder model is first trained on event data ($\pazocal{D}_{S}$) for the next activity prediction, as no extra labels are required for this target. The last hidden state of the encoder model’s final LSTM layer is then transferred to a second model for outcome prediction. This model is then fine-tuned on a portion of the same event data, for which a process-related outcome label is available ($\pazocal{D}_{T}$).
@@
Besides parameter-based TL, Cabrera et al. (2022) propose a PPM technique that leverages feature-based TL. This is achieved by using various BERT models – pre-trained on a huge corpus of text data ($\pazocal{D}_{S}$) – to encode the text information available in event data ($\pazocal{D}_{T}$) and adding the encoded information to the input of a multi-tasking LSTM model. This LSTM model addresses the next activity prediction and the next timestamp prediction. In addition, the work suggests that fine-tuning the pre-trained BERT models on available event data using the next activity prediction task enhances the LSTM model’s performance for predicting next activities and timestamps as it adapts the representation of the BERT models to the process control-flow.
@@
Some PPM techniques in the first group also perform both parameter- and feature-based TL. Pfeiffer et al. (2021) propose a PPM technique combining Gramian angular fields and representation learning. The Gramian angular fields are applied to encode the event data as 2D images, while representation learning is performed to obtain a generic process representation from the encoded event data ($\pazocal{D}_{S}$). This representation is obtained using a two-step approach. First, a CNN model is trained to map the encoded input onto trace variants. Second, the pre-trained CNN model is extended with a fully-connected layer and heads for various prediction tasks, such as the next activity and next timestamp prediction, and then fine-tuned on the encoded event data. Subsequently, all heads except the desired one are removed, and the model is fine-tuned again on the encoded input ($\pazocal{D}_{T}$). By learning a model capturing a generic process representation and aligning it to various prediction tasks, feature- and parameter-based TL is performed within one model. In a similar approach, Chen et al. (2022) propose to learn a generic representation of traces by pre-training a BERT model on a set of unlabeled trace data ($\pazocal{D}_{S}$) using a custom masked activity modeling task. The pre-trained BERT model is then fine-tuned on another set of labeled trace data ($\pazocal{D}_{S}$) for the next activity or outcome prediction. As in the approach of Pfeiffer et al. (2021), one model capturing a generic process representation is learned and aligned to prediction tasks. Therefore, this approach also performs feature- and parameter-based TL. Pasquadibisceglie et al. (2023) suggest using TL for handling concept drifts in PPM. Specifically, once a drift is detected within event data ($\pazocal{D}_{S}$), the parameters of both an LSTM model for the next activity prediction and a word2vec model for activity encoding are updated with current event data ($\pazocal{D}_{T}$). This enhances the prediction performance of the LSTM model and enables the word2vec model to deal with newly occurring activities. However, even though these PPM techniques incorporate parameter- and feature-based TL, they are not designed to work with different business processes, where the values and distributions of event attributes in the source and target domains are generally different. Consequently, these PPM techniques cannot be used in an intra- or inter-organizational TL setting.
@@
Pfeiffer et al. (2021) Inductive, homogeneous Parameter-based, feature-based CNN model with task-specific layer ACT, TS, RES, DA Transfer process knowledge from a model capturing a generic business process representation for various single prediction tasks to prediction models for single tasks, to address them more effectively Integer encoding, min-max scaling, Gramian angular field transformation Extraction of one time-related feature (duration since start), min-max normalization, Gramian angular field transformation RTP, NAP, OP Helpdesk, BPIC12, BPIC13, BPIC17, BPIC20, MobIS
Chen et al. (2022) Inductive, homogeneous Parameter-based, feature-based Multi-tasking, BERT-based transformer model ACT Transfer process knowledge from a generic business process representation for multiple tasks to prediction models for single tasks, to address them more effectively Embedding layer, positional encoding - NAP, OP Helpdesk, BPIC12, Sepsis
@@
Cao et al. (2022) Transductive, homogeneous Parameter-based MLP model ACT, TS Transfer process knowledge from a prediction model pre-trained on event data (previous transition division) to another prediction model fine-tuned on event data (current transition division), to address a prediction task more effectively One-hot encoding, autoencoder Extraction of eight time-related features (e.g., month, execution time, and elapsed time), min-max normalization, autoencoder RTP Four synthetic event logs, BPIC12, BPIC17,
@@
Cabrera et al. (2022) Inductive, heterogeneous Feature-based Multi-tasking LSTM model ACT, TS, text Transfer process knowledge from a generic embedding model pre-trained on a large corpus of text data to an embedding model fine-tuned on real-life event data, to address prediction tasks with a successive prediction model more effectively Embedding model Extraction of three time-related features (time passed since previous event, time within the day, and time within the week), min-max normalization NAP, NTP BPIC16
@@
Nur et al. (2024) Transductive, heterogeneous Parameter-based, feature-based Transformer model with time-related feature block ACT, TS Transfer process knowledge from a prediction model pre-trained on real-life event data to another prediction model fine-tuned on real-life event data, to address a prediction task more effectively Embedding layer, positional encoding Extraction of two time-related features (elapsed time and lagged time) RTP Helpdesk, BPIC20, BPIC12, BPIC11
@@
Furthermore, two PPM techniques combine feature-based and parameter-based TL. Rizk et al. (2023) pre-train an LSTM model on synthetic event data ($\pazocal{D}_{S}$), transfer parameters of this model to a new one, and fine-tune the new model using other synthetic event data ($\pazocal{D}_{T}$) to solve the same prediction task (i.e., next activity prediction). The first layer of the LSTM model is an embedding layer, which enables matching the feature space of $\pazocal{D}_{S}$ to that of $\pazocal{D}_{T}$. The same experiment is performed for the event data of two different real-life event logs. Nur et al. (2024) suggest a technique with a transformer-based DL model for remaining time prediction. The transformer-based DL model consists of three components: a trace and positional encoding block, a transferable block (leveraging multi-head self-attention to capture semantic and temporal relationships), and a time-related feature block. TL is used in the technique to address the challenge of limited event data. This involves training the transformer-based DL model (base model) on a large amount of event data ($\pazocal{D}_{S}$), transferring the trained parameters of its transferable block to a new model, and fine-tuning the new model on a smaller amount of event data ($\pazocal{D}_{T}$) to improve its prediction performance for remaining process times. Although both techniques incorporate the ideas of feature-based and parameter-based TL in their models, they require a fine-tuning step on event data in the target context. However, this is difficult in intra- or inter-organizational TL settings, where event data are not or very limitedly available in the target context.
Table 1 summarizes the works identified, which use TL for PPM, and shows how they differ from our proposed technique in terms of dimensions, such as transfer goal, different business processes, and fine-tuning in target not necessary.
@@ 5 Transfer learning-based technique for predictive process monitoring
This section describes the design of our proposed TL-based technique for PPM. The technique is structured into three phases, as shown in Figure 11. In the first phase, initial model building on source, an event log from the source context is loaded, preprocessed, and a DNN-based base model is built and evaluated on preprocessed event data. In the second phase, transfer resources from source to target, relevant resources from the source context (e.g., pre-trained models) are transferred to the target context. If an event log from the target context is available, the base model from the source context can be applied and evaluated in an intermediate step before being used for (ongoing) process instances in the target context. In the third phase, online application of model to target, an ongoing process instance of the target context is received. This process instance is then pre-processed using the transferred resources (e.g., information for timestamp encoding), and the transferred base model is applied to the pre-processed data. Finally, operational support is provided based on the base model’s predictions for process improvement. In the following sections, the three phases of our proposed TL-based technique for PPM are described.